OpenStudy (johnt):

Does ln(sec(x)) = - ln(cos(x)) ? The graphs look the same, but if I gave this answer on a test, would it be accurate?

OpenStudy (johnt):

\[\ln \left( \sec x \right) = - \ln \left( \cos x \right)\]

hartnn (hartnn):

yes, it is same \(\ln (1/a)= \ln (a^{-1})= -\ln a \) here, a = cos x = 1/ sec x, right ?

OpenStudy (johnt):

Ah, I see. Your equation with the power example makes sense.

hartnn (hartnn):

to be more accurate, i'll suggest you to write : \(-\ln|\cos x| \) because log of negative numbers is not defined...